OpenStudy (anonymous):

What is the solution to m+ (-10)=-33

OpenStudy (wach):

Let's simplify this: m - 10 = -33

OpenStudy (wach):

If we want to balance this expression to get the value for m, we would want to add 10 to both sides.

OpenStudy (anonymous):

m=-23

OpenStudy (charlotte123):

m+ (-10) = -33 + 10 + 10 m = -23

OpenStudy (charlotte123):

@zamani12345 Correct! :)

OpenStudy (wach):

Wow yeah you're right, that was a really stupid mistake >_< -33 +10 = -23